OpenStudy (anonymous):

Sizes of envelopes. Rhetoric Advertising is a direct-mail company. It determines that for a particular campaign, it can use any envelope with a fixed width of 3 in. and an area of at least 17 in2. Determine (in terms of an inequality) those lengths that will satisfy the company constraints.@satellite73

OpenStudy (anonymous):

\[3x\geq 17\] \[x\geq\frac{17}{3}\]

OpenStudy (anonymous):

i made a mistake its 3 1/2 &17 1/2 in^2